factors for
a)6xsquared-5x-5+0
b)3xsquared-7x+3
c) xcubed+2xsquared-3x

Answer by stanbon(75887)

factors for
a)6x^2-5x-5
Use the quadratic formula to get:
x=[5+sqrt(145)]/12 or [5-sqrt(145)]/12
---------------------------------------
b)3x^2-7x+3
Use the quadratic formula to get:
x=[7+sqrt(-36)]/6 or x=[7-sqrt(-36)]/6
x=[7+6i]/6 or x=[7-6i]/6
--------------------------------
c) x^3+2x^2-3x
=x(x^2+2x-3)
=x(x+3)(x-1)
-----------------------
Cheers,
Stan H.